OpenStudy (awkwardpanda):

find the sum of the following series. 15 ∑ n=1 (2n+1) a.240 b.255 c.210 d.510

OpenStudy (anonymous):

there is some sort of formula for this, but i don't know it on the other hand \[\sum (2n+1)=2\sum n +\sum 1\] and \[\sum n=\frac{n(n+1)}{2}\] also \[\sum 1=n\]

OpenStudy (anonymous):

that gives \[2\sum (n+1)=2\times \frac{n(n+1)}{2}+n=n(n+1)+n=n^2+2n\]

OpenStudy (awkwardpanda):

you're my favorite person.

OpenStudy (anonymous):

(blush)